# InfFeed: Influence Functions as a Feedback to Improve the Performance of Subjective Tasks

### 22 Feb 2024 v1

@@ 3 Preliminaries
$IF\{(x_{i},y_{i}),(x_{ts},y_{ts})\}\cong\frac{d\mathcal{L}((x_{ts},y_{ts}),\hat{\theta}_{\pm{\delta},(x_{i},y_{i})})}{d({\pm{\delta}})}$ (1)
@@
$\hat{\theta}={\arg\min}_{\theta}\frac{1}{n}\sum_{i=1}^{n}\{\mathcal{L}((x_{ts},y_{ts}),\theta)+({\pm{\delta}})\mathcal{L}((x_{i},y_{i}),\theta)\}$ (2)
@@
$\begin{split}&IF\{(x_{i},y_{i}),(x_{ts},y_{ts})\}\approx\\ &-\nabla_{\theta}\mathcal{L}((x_{ts},y_{ts}),\hat{\theta})^{T}{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\mathcal{L}((x_{i},y_{i}),\hat{\theta})\end{split}$ (3)
@@ 4.1 Influence function to introduce feedback
System 2 (InfFeed): We begin by partitioning the training set [^3] https://docs.cleanlab.ai/v2.0.0/tutorials
/pred_probs_cross_val.html
, denoted by $T_{R}$, into a smaller subset $T_{CR}$, which we designate as a fine-tuning set. Using the remaining part of the training set, $T_{PR}=T_{R}-T_{CR}$, we then train a model, $\theta_{A}$. For each instance in $T_{CR}$, we determine the most influential training instances from $T_{PR}$, with $\theta_{A}$ and the influence function approach outlined in the preceding section. We revise the label of each instance in $T_{CR}$ based on the majority/weighted voting of the labels from the top-$K$ influential instances identified earlier, producing an updated set $T_{CR}^{up}$. We proceed to fine-tune $\theta_{A}$ using $T_{CR}^{up}$. Afterwards, we utilize the held-out validation set $V$ to derive the final model, $\theta_{B}$. Finally, we evaluate $\theta_{B}$ using the held-out test dataset $T_{S}$ (Figure 2, System 2).
@@
Baselines: In this paper, we use four state-of-the-art baseline methods taken from the literature – Hao et al. Hao et al. (2020), Rajani et al. Rajani et al. (2020), Wang et al. Wang et al. (2020), and Kong et al. Kong et al. (2022). As additional baselines, we use two state-of-the-art LLMs GPT-3.5-Turbo [^4] https://platform.openai.com/docs/models/gpt-3-5 and GPT-4555https://platform.openai.com/docs/models/gpt-4, in a zero-shot classification setting.

### 09 Mar 2024 v2

@@ 3 Preliminaries
$\footnotesize IF\{(x_{i},y_{i}),(x_{ts},y_{ts})\}\cong\frac{d\mathcal{L}((x_{ts},y_{ts}),\hat{\theta}_{\pm{\delta},(x_{i},y_{i})})}{d({\pm{\delta}})}$ (1)
@@
$\footnotesize\hat{\theta}={\arg\min}_{\theta}\frac{1}{n}\sum_{i=1}^{n}\{\mathcal{L}((x_{ts},y_{ts}),\theta)+({\pm{\delta}})\mathcal{L}((x_{i},y_{i}),\theta)\}$ (2)
@@
$\footnotesize\begin{split}&IF\{(x_{i},y_{i}),(x_{ts},y_{ts})\}\approx-\nabla_{\theta}\mathcal{L}((x_{ts},y_{ts}),\hat{\theta})^{T}{H}_{\hat{\theta}}^{-1}\nabla_{\theta}\mathcal{L}((x_{i},y_{i}),\hat{\theta})\end{split}$ (3)
@@ 4.1 Influence function to introduce feedback
System 2 (InfFeed): We begin by partitioning the training set [^3] https://docs.cleanlab.ai/v2.0.0/tutorials/pred_probs_cross_val.html, denoted by $T_{R}$, into a smaller subset $T_{CR}$, which we designate as a fine-tuning set. Using the remaining part of the training set, $T_{PR}=T_{R}-T_{CR}$, we then train a model, $\theta_{A}$. For each instance in $T_{CR}$, we determine the most influential training instances from $T_{PR}$, with $\theta_{A}$ and the influence function approach outlined in the preceding section. We revise the label of each instance in $T_{CR}$ based on the majority/weighted voting of the labels from the top-$K$ influential instances identified earlier, producing an updated set $T_{CR}^{up}$. We proceed to fine-tune $\theta_{A}$ using $T_{CR}^{up}$. Afterwards, we utilize the held-out validation set $V$ to derive the final model, $\theta_{B}$. Finally, we evaluate $\theta_{B}$ using the held-out test dataset $T_{S}$ (Figure 2, System 2).
@@
Baselines: In this paper, we use four state-of-the-art baseline methods taken from the literature - Hao et al. Hao et al. (2020), Rajani et al. Rajani et al. (2020), Wang et al. Wang et al. (2020), and Kong et al. Kong et al. (2022). As additional baselines, we use two state-of-the-art LLMs GPT-3.5-Turbo [^4] https://platform.openai.com/docs/models/gpt-3-5 and GPT-4555https://platform.openai.com/docs/models/gpt-4, in a zero-shot classification setting.